OpenStudy (anonymous):

how to find the answer to the following: find two consecutive integers whose sum is 45. show work plz

OpenStudy (ccswims):

22 and 23 22 + 23 = 45

OpenStudy (ccswims):

and 22 and 23 are consecutive

OpenStudy (isaiah.feynman):

Explain from scratch how you got there.

OpenStudy (ccswims):

Ok, well division won't help in this case, and I know 25 + 25 = 50 so I just need to subtract 5 from both of the numbers put together ( 2 + 3 = 5 , and then add the 20 is there, 22 + 23 = 25)

OpenStudy (isaiah.feynman):

Well a good way to show is by saying this..let those numbers be n and n+1

OpenStudy (isaiah.feynman):